OpenStudy (anonymous):

PLEASE HELP!!! Select the equation of a line that is parallel to the line on the graph and passes through the point (2, 3).

OpenStudy (anonymous):

a.y = 2x + 4 b. y = - 2x + 2 c.y = - x + 2 d.y = - x + 4

OpenStudy (anonymous):

@zepdrix @BVBLegacy @Compassionate please help

zepdrix (zepdrix):

Hey Nicki c: Do you know how to find the `slope` of the line that is graphed?

OpenStudy (anonymous):

Slope = rise over run right?

zepdrix (zepdrix):

Yes, so you'll want to use two points to find the rise over run. It might be easier if you use the intercepts.

zepdrix (zepdrix):

(-2,0) and (0,-1), yah?

OpenStudy (anonymous):

Yea

OpenStudy (anonymous):

But then what do i do next ? :/ @zepdrix .. Sorry im really bad at math lol.

zepdrix (zepdrix):

So your slope formula is given by:\[\Large\rm m=\frac{y_2-y_1}{x_2-x_1}\]On the bottom of the equation we're finding the space between the two points (run), and on the top we're finding the height between them (rise). Understand how to plug that information in? :o Find your slope! c: It'll be a few more steps after that, but let's get this figured out first.

OpenStudy (anonymous):

\[m= \frac{ -1-0 }{ -2 - 0 } \] ?

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

Woops you plugged one of the points in backwards D:\[\Large\rm (-2,0)\qquad (0,-1)\] \[\Large\rm m=\frac{0-(-1)}{-2-0}\]

OpenStudy (anonymous):

oops . So after that what do i do .-. ?

zepdrix (zepdrix):

After simplifying, we determine that the line that is graphed has a slope of -1/2. We're trying to construct a line which is parallel to that one. Remember what's special about parallel lines? They have the same slope.

zepdrix (zepdrix):

Here are two of the ways we can write a line, Slope-Intercept Form:\[\Large\rm y=mx+b\] Point-Slope Form:\[\Large\rm y-y_1=m(x-x_1)\] Have you seen the second equation before? It's easier to use, less work for you. But if you're not familiar with it, we can stick with the slope-intercept form.

OpenStudy (anonymous):

ok, im not really familiar with it. I dont really understand slope or any of that..... @zepdrix

zepdrix (zepdrix):

You're trying to find a line that is parallel to the one pictured. We found the slope of the line pictured. Since they're parallel, we know that our new line will have the same slope value. \(\Large\rm y=-\frac{1}{2}x+b

zepdrix (zepdrix):

\(\Large\rm y=-\frac{1}{2}x+b\)

zepdrix (zepdrix):

To find the missing y-intercept, plug in the point that they gave you. \(\Large\rm (2,3)\)

zepdrix (zepdrix):

\[\Large\rm 3=-\frac{1}{2}\cdot 2+b\]

zepdrix (zepdrix):

Solve for b, then plug it back into here,\[\Large\rm y=-\frac{1}{2}x+b\]

OpenStudy (anonymous):

ok, so b = 4? @zepdrix

OpenStudy (anonymous):

@mathmath333 please help <

OpenStudy (mathmath333):

\(\huge b=4\)

OpenStudy (anonymous):

thank you ! @mathmath333

OpenStudy (anonymous):

nickiconda is right shes awesome